(800/8*3+25)-(20+20*6/2)
1
Expert's answer
2010-07-08T03:51:01-0400
(800/8*3+25)-(20+20*6/2) = ((100*3)+25)-(20+(20*3)) = (300+25)-(20+60) = 325 - 80 = 245